OpenStudy (anonymous):

consider a saving account which pays a nominal interest at 8% per annum, paid quarterly. Calculate the effective annual interest rate per annum

OpenStudy (anonymous):

ive looked for it on the internet to be honest i found something that might help : r = (1+ i/n)^n - 1 r - effective annual rate i - nominal rate n - number of compounding periods per year r = (1 + 0.08/4)^4 - 1 = 0.0824 -> 8.24% http://en.wikipedia.org/wiki/Effective_interest_rate hope it helps

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you are welcome